### v1

@@ 3 Examples
The following examples illustrate the notion of BCWE in two basic games and show how to implement the social cost-minimizing BCWE as an obedient BWE of the basic game extended with a direct information structure. They also show that restricting attention to public information or deterministic flows is not without loss of generality.
Example 1.
[…]
$c_{a}(\boldsymbol{y},\theta)=\operatorname{\mathds{1}}\{\theta=\bar{\theta}\},\quad c_{b}(\boldsymbol{y},\theta)=2y_{b}+1/3.$ (3.1)
Assume that the designer aims at minimizing the social cost, defined as
$\operatorname{\mathsf{SC}}(\boldsymbol{y},\theta)=y_{a}c_{a}(\boldsymbol{y},\theta)+y_{b}c_{b}(\boldsymbol{y},\theta).$
[…]
This outcome can be implemented as a BWE with the following information structure: there are six populations of size $\gamma^{k}=1/6$ for $k=1,\ldots,6$, with
[…]
and
$\pi(b,a,a,a,a,a\mid\bar{\theta})=\pi(a,b,a,a,a,a\mid\bar{\theta})=\pi(a,a,b,a,a,a\mid\bar{\theta})=$
$\pi(a,a,a,b,a,a\mid\bar{\theta})=\pi(a,a,a,a,b,a\mid\bar{\theta})=\pi(a,a,a,a,a,b\mid\bar{\theta})=$ $1/6$
It is immediate to observe that the equilibrium conditions of the BWE are equivalent to those of the corresponding BCWE.
Example 2.
In the following basic game from Díaz et al., (2009) and Mitsche et al., (2013), there is a single state, and the action set is $\mathcal{A}=\left\{a,b\right\}$. The cost functions are:
$c_{a}(\boldsymbol{y})=1,\quad c_{b}(\boldsymbol{y})=\max\left\{2-4y_{b},4y_{b}-2\right\}.$ (3.3)
Action $b$ generates negative cost externalities when $y_{b}<1/2$ and positive cost externalities when $y_{b}>1/2$. This basic game without information admits three BWE outcomes: $(1,0)$, $(3/4,1/4)$, and $(1/4,3/4)$, all of which yield a social cost of $1$. Public information just induces a mixing over these outcomes. Now consider the flows $\boldsymbol{z}=(1,0)$ and $\boldsymbol{w}=(1/2,1/2)$. The outcome $\mu$, defined by
@@
$\frac{2}{3}=\frac{1}{3}z_{a}c_{a}(\boldsymbol{z})+\frac{2}{3}w_{a}c_{a}(\boldsymbol{w})$ $\leq\frac{1}{3}z_{a}c_{b}(\boldsymbol{z})+\frac{2}{3}w_{a}c_{b}(\boldsymbol{w})=\frac{2}{3},$
$0=\frac{1}{3}z_{b}c_{b}(\boldsymbol{z})+\frac{2}{3}w_{b}c_{b}(\boldsymbol{w})$ $\leq\frac{1}{3}z_{b}c_{a}(\boldsymbol{z})+\frac{2}{3}w_{b}c_{a}(\boldsymbol{w})=\frac{1}{3}.$
The social cost of this BCWE is $2/3$, which is lower than the social cost with public information. The cost-minimizing BCWE is (partially) implemented as a BWE with the following direct information structure: there are two populations of size $\gamma^{1}=\gamma^{2}=1/2$, and
$\pi(a,a)=\pi(a,b)=\pi(b,a)=\frac{1}{3}.$
@@ 4 Partial Implementation
It is immediate to show that the set of all BWE for all possible (direct or indirect) information structures, as described in the previous section, is included in the set of BCWE. This simple observation explicitly illustrates how to construct some BCWE: choose an information structure $\left(\boldsymbol{\gamma},\mathcal{T},\pi\right)$ used to disclose information, and choose a BWE of the basic game extended with this information structure.
The next theorem establishes, at least approximately, the other direction of the revelation principle (Myerson,, 1982, Bergemann and Morris,, 2016). Specifically, it shows that if $\mu$ is a BCWE with finite support and rational flows, then there exists a direct information structure such that $\mu$ is implemented as an obedient BWE outcome. If the support of $\mu$ is not finite or if some flows in the support are not rational, then, for every $\varepsilon>0$, $\mu$ can be approximately implemented as a Bayesian Wardrop $\varepsilon$-equilibrium outcome.
@@
Let $\mu$ be a BCWE and $\psi\colon\mathcal{Y}\times\Theta\to\mathbb{R}$ be a continuous function. For all $\varepsilon>0$, there exists a direct information structure $\mathcal{I}$ such that the obedient interim flow profile $\breve{\boldsymbol{y}}_{\mathcal{I}}$ is a Bayesian Wardrop $\varepsilon$-equilibrium of $(\Gamma,\mathcal{I})$ and,
@@
In Theorem 4.1, $\psi$ represents any possible cost function for the designer, and (4.1) means that, for every state, the interim expected cost of the designer under outcome $\breve{\mu}_{\mathcal{I}}$ is $\varepsilon$-close to the interim expected cost under the BCWE $\mu$. The proof of the theorem also shows that, for every $L>0$, the same $\mathcal{I}$ can be chosen for all $L$-Lipschitz $\psi$.
Proof.
Koessler et al., (2024, Lemma 2) show that BCWE with finite support are dense within the set of Bayes correlated Wardrop equilibria. Therefore, it is enough to prove the result for $\mu$ with finite support. Fix $\varepsilon>0$ and a BCWE $\mu$ with finite support $\mathcal{Y}^{\ast}\subset\mathcal{Y}$. Let
$\varepsilon_{0}=\min\left\{y_{a}\text{ s.t. }\boldsymbol{y}\in\mathcal{Y}^{\ast},a\in\mathcal{A},y_{a}>0\right\}>0$ (4.2)
be the smallest positive flow in the finite set $\mathcal{Y}^{\ast}$.
We take an arbitrary $\eta>0$ and approximate flows in $\mathcal{Y}^{\ast}$ with vectors having rational components as follows. There exists $K\in\mathbb{N}_{+}$, such that for every $\boldsymbol{y}\in\mathcal{Y}^{\ast}$ and every $a\in\mathcal{A}$, there exists $N(y_{a})\in\mathbb{N}$, such that,
$\sum_{a\in\mathcal{A}}N(y_{a})=K\quad\text{and}\quad\forall a\in\mathcal{A},\left\lvert\frac{N(y_{a})}{K}-y_{a}\right\rvert\leq\eta.$ (4.3)
For each $\boldsymbol{y}\in\mathcal{Y}^{\ast}$, let $\tilde{\boldsymbol{y}}$ denote the flow profile $(\tilde{y}_{a})_{a\in\mathcal{A}}$ with $\tilde{y}_{a}=N(y_{a})/K$.
We construct a direct information structure $\mathcal{I}$ as follows. There are $|\mathcal{K}|=K$ populations and each population $k$ has the same size $\gamma^{k}=1/K$. For each population $k$, the set of types is $\mathcal{T}^{k}=\mathcal{A}$. Conditionally on state $\theta$, a flow profile $\boldsymbol{y}\in\mathcal{Y}^{\ast}$ is drawn with probability $\mu(\boldsymbol{y}\mid\theta)$, then action $a$ is recommended to a subset of $N(y_{a})$ populations chosen uniformly, so that, conditionally on $\boldsymbol{y}$ and $\theta$, the probability that population $k$ is recommended $a$ is $\tilde{y}_{a}$. We let $\breve{\mu}_{\mathcal{I}}$ be the induced obedient outcome which puts probability $\mu(\boldsymbol{y}\mid\theta)$ on $\tilde{\boldsymbol{y}}$ in state $\theta$.
For each population $k$, the conditional expected cost of playing $b$ when $a$ is recommended is
$\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}\frac{p(\theta)\mu(\boldsymbol{y}\mid\theta)\tilde{y}_{a}}{\operatorname{\mathsf{Q}}(a)}c_{b}(\tilde{\boldsymbol{y}},\theta),$ (4.4)
@@
$\operatorname{\mathsf{Q}}(a)=\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}p(\theta)\mu(\boldsymbol{y}\mid\theta)\tilde{y}_{a}.$ (4.5)
For each pair of actions $(a,b)$, the mapping $\boldsymbol{y}\mapsto y_{a}c_{b}(\boldsymbol{y},\theta)$ is uniformly continuous on the compact set $\mathcal{Y}$. Thus, there exists a common modulus of continuity $\omega(\,\cdot\,)$ with $\lim_{\eta\searrow 0}\omega(\eta)=0$ such that, for all $\theta\in\Theta$, for all $(a,b)\in\mathcal{A}\times\mathcal{A}$, for all $(\boldsymbol{y},\boldsymbol{z})\in\mathcal{Y}\times\mathcal{Y}$,
$\left\lvert y_{a}c_{b}(\boldsymbol{y},\theta)-z_{a}c_{b}(\boldsymbol{z},\theta)\right\rvert\leq\omega(\max_{b}\lvert y_{b}-z_{b}\rvert).$ (4.6)
@@
$\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}p(\theta)\mu(\boldsymbol{y}\mid\theta)y_{a}c_{a}(\boldsymbol{y},\theta)\leq\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}p(\theta)\mu(\boldsymbol{y}\mid\theta)y_{a}c_{b}(\boldsymbol{y},\theta).$ (4.7)
It follows from Eqs. 4.7, 4.3 and 4.6 that
$\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}p(\theta)\mu(\boldsymbol{y}\mid\theta)\tilde{y}_{a}c_{a}(\tilde{\boldsymbol{y}},\theta)\leq\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}p(\theta)\mu(\boldsymbol{y}\mid\theta)\tilde{y}_{a}c_{b}(\tilde{\boldsymbol{y}},\theta)+2\omega(\eta),$ (4.8)
and thus
$\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}\frac{p(\theta)\mu(\boldsymbol{y}\mid\theta)\tilde{y}_{a}}{\operatorname{\mathsf{Q}}(a)}c_{a}(\tilde{\boldsymbol{y}},\theta)\leq\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}\frac{p(\theta)\mu(\boldsymbol{y}\mid\theta)\tilde{y}_{a}}{\operatorname{\mathsf{Q}}(a)}c_{b}(\tilde{\boldsymbol{y}},\theta)+\frac{2\omega(\eta)}{\operatorname{\mathsf{Q}}(a)}.$ (4.9)
From Eq. 4.3 we have
$\operatorname{\mathsf{Q}}(a)=\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}p(\theta)\mu(\boldsymbol{y}\mid\theta)\tilde{y}_{a}\geq\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}p(\theta)\mu(\boldsymbol{y}\mid\theta)(y_{a}-\eta)\geq\varepsilon_{0}-\eta\geq\frac{\varepsilon_{0}}{2},$ (4.10)
@@
Now, take a continuous $\psi\colon\mathcal{Y}\times\Theta\to\mathbb{R}$. Since $\mathcal{Y}\times\Theta$ is compact, $\psi$ is uniformly continuous and admits a modulus of continuity $\omega_{\psi}$ such that for all $\theta\in\Theta$, for all $(a,b)\in\mathcal{A}\times\mathcal{A}$, for all $(\boldsymbol{y},\boldsymbol{z})\in\mathcal{Y}\times\mathcal{Y}$,
$\left\lvert\psi(\boldsymbol{y},\theta)-\psi(\boldsymbol{z},\theta)\right\rvert\leq\omega_{\psi}(\max_{b}\lvert y_{b}-z_{b}\rvert).$ (4.12)
We have for each $\theta\in\Theta$,
$\begin{split}\left\lvert\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}\mu(\boldsymbol{y}\mid\theta)\psi(\tilde{\boldsymbol{y}},\theta)-\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}\mu(\boldsymbol{y}\mid\theta)\psi(\boldsymbol{y},\theta)\right\rvert&\leq\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}\mu(\boldsymbol{y}\mid\theta)\left\lvert\psi(\tilde{\boldsymbol{y}},\theta)-\psi(\boldsymbol{y},\theta)\right\rvert\\ &\leq\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}\mu(\boldsymbol{y}\mid\theta)\omega_{\psi}(\eta)\\ &\leq\omega_{\psi}(\eta),\end{split}$ (4.13)
@@
If flows in $\mathcal{Y}^{\ast}$ have rational coefficients with common denominator $K$, then we can choose $\tilde{\boldsymbol{y}}=\boldsymbol{y}$ and $\eta=0$, we have a Bayesian Wardrop 00-equilibrium. Finally, notice that if $\psi$ is $L$-Lipschitz (for the sup norm on flows), then we can choose $\omega_{\psi}(\eta)=L\eta$. Thus, for given $\varepsilon$, the choice of $\eta$ and $\mathcal{I}$ is the same across all such $\psi$. ∎
5 Full Implementation
In this section, we study full implementation of BCWE in games with convex potential.

### v2

@@ 3 Examples
The following examples illustrate the notion of BCWE in two simple games and show how to implement the total welfare-maximizing BCWE as an obedient BWE of the game extended with a direct information structure. The examples also show that restricting attention to public information or deterministic flows is not without loss of generality in terms of efficiency.
Example 3.1 (Platform access with network effects and congestion).
In the following game, studied by Mitsche et al., (2013), there is a single state, and the action set is $\mathcal{A}=\left\{a,b\right\}$. The payoff functions are:
$u_{a}(\boldsymbol{y})=1,\quad u_{b}(\boldsymbol{y})=\min\left\{4y_{b},4-4y_{b}\right\}.$ (3.1)
Eq.˜3.1 says that action $b$ generates positive payoff externalities when $y_{b}<1/2$ and negative payoff externalities when $y_{b}>1/2$. This can be interpreted as a platform access decision, where action $b$ corresponds to joining a platform whose value increases with moderate participation due to network effects but deteriorates when participation becomes too high because of congestion, while action $a$ represents a safe outside option with constant payoff 1. This game without information admits three BWE outcomes: $(1,0)$, $(3/4,1/4)$, and $(1/4,3/4)$, all of which yield a total payoff of $1$. Public information induces a mixing over these outcomes, so the expected total payoff remains $1$. Consider now the flows $\boldsymbol{z}=(1,0)$ and $\boldsymbol{w}=(1/2,1/2)$. The outcome $\mu$, defined by
@@
$\begin{split}\frac{2}{3}=\frac{1}{3}z_{a}u_{a}(\boldsymbol{z})+\frac{2}{3}w_{a}u_{a}(\boldsymbol{w})&\geq\frac{1}{3}z_{a}u_{b}(\boldsymbol{z})+\frac{2}{3}w_{a}u_{b}(\boldsymbol{w})=\frac{2}{3},\\ \frac{2}{3}=\frac{1}{3}z_{b}u_{b}(\boldsymbol{z})+\frac{2}{3}w_{b}u_{b}(\boldsymbol{w})&\geq\frac{1}{3}z_{b}u_{a}(\boldsymbol{z})+\frac{2}{3}w_{b}u_{a}(\boldsymbol{w})=\frac{1}{3}.\end{split}$ (3.3)
The total payoff of this BCWE is $4/3$, which is higher than the total payoff with public information. From the platform’s perspective, private recommendations are used to correlate users’ decisions so that the first-best participation level $(1/2,1/2)$ is achieved with probability $2/3$, while the inefficient outcome in which all players choose $b$ never occurs.
The welfare-maximizing BCWE is (partially) implemented as a BWE with the following direct information structure: there are two populations of size $\gamma^{1}=\gamma^{2}=1/2$, and
$\pi(a,a)=\pi(a,b)=\pi(b,a)=\frac{1}{3}.$ (3.4)
Operationally, the platform can achieve this outcome by partitioning users into two equally sized groups and privately recommending access to at most one group at a time.
Example 3.2 (Entry on a market with uncertain production cost).
We consider a game with binary action set $\mathcal{A}=\left\{a,b\right\}$, where action $a$ indicates staying out of the market, and action $b$ indicates entering the market. Firms on the market produce one unit each and sell it at a market price decreasing with total production. The state $\theta$ is the common production cost, which can take values 0 or $1$. This game is a particular instance of the Cournot model described in Section˜5.1.
So we let $\Theta=\left\{0,1\right\}$, and for $q\in(0,1)$,
$p(0)=1-q,\quad\text{and}\quad p(1)=q.$ (3.5)
Let us consider deterministic outcomes, where the flow profile in state $\theta$ is some $\boldsymbol{y}(\theta)=\left(y_{a}(\theta),y_{b}(\theta)\right)=\left(1-y_{b}(\theta),y_{b}(\theta)\right)\in\mathcal{Y}$. The inverse demand depends linearly on total production and is given by
$\rho(y_{b}(\theta))=1-y_{b}(\theta).$ (3.6)
Firms maximize profits, which are
$\begin{split}u_{a}(\boldsymbol{y},\theta)&=0,\\ u_{b}(\boldsymbol{y},\theta)&=\rho(y_{b}(\theta))-\theta.\end{split}$ (3.7)
The total profit in state $\theta$ is
$y_{a}(\theta)u_{a}(\boldsymbol{y}(\theta),\theta)+y_{b}(\theta)u_{b}(\boldsymbol{y}(\theta),\theta)=y_{b}(\theta)(1-y_{b}(\theta)-\theta).$ (3.8)
The expected total profit is
$(1-q)\,y_{b}(0)(1-y_{b}(0))-q\,y_{b}(1)y_{b}(1).$ (3.9)
The maximizer over $0,1]^{2}$ is $\left(y_{b}(0)=1/2,y_{b}(1)=0\right)$, with expected total profit equal to (1q)/4(1-q)/4.
For deterministic outcomes, the BCWE equilibrium constraints reduce to:
$(1-q)(1-y_{b}(0))^{2}$ $\leq q\,y_{b}(1)(1-y_{b}(1)),$ (3.10)
$(1-q)y_{b}(0)(1-y_{b}(0))$ $\geq q\,(y_{b}(1))^{2}.$ (3.11)
Maximizing expected total profit subject to Eqs.˜3.11 and 3.10 yields a unique solution with (3.10) binding and (3.11) slack:
$y_{b}^{*}(0)=1-\frac{\sqrt{q}}{2},\quad y_{b}^{*}(1)=\frac{1-\sqrt{q}}{2}.$ (3.12)
The associated expected total profit is $\left(\sqrt{q}-q\right)/2$, which is strictly smaller than the unconstrained benchmark but strictly larger than the expected total profit achievable under public information, i.e., 0. By Proposition˜6.1, restricting attention to deterministic flows is without loss of optimality in this binary-action environment, so $(y_{b}^{*}(0),y_{b}^{*}(1))$ is the total profit-maximizing BCWE outcome.
Similarly to the previous example, as long as $\sqrt{q}$ is a rational number, this deterministic BCWE can be implemented as an obedient BWE using a direct information structure with finitely many populations, by recommending action $b$ to a fraction $y_{b}^{*}(\theta)$ of firms in each state $\theta$. For example, if $q=1/4$, then
$y_{b}^{*}(0)=1-\frac{\sqrt{q}}{2}=\frac{3}{4},\qquad y_{b}^{*}(1)=\frac{1-\sqrt{q}}{2}=\frac{1}{4}.$ (3.13)
Consider four populations of equal size $\gamma^{k}=1/4$ for every $k$, and a direct information structure such that, in state 0, exactly three populations are recommended action $b$ and one population is recommended action $a$, whereas in state $1$ exactly one population is recommended action $b$ and three populations are recommended action $a$ (uniformly over the choice of the recommended populations). This induces the deterministic outcome $(y_{b}(0),y_{b}(1))=(3/4,1/4)$. As shown in Section˜5, this total profit-maximizing outcome is fully implemented in this game because it is a concave potential game. Moreover, when $\sqrt{q}$ is not rational, the total profit-maximizing outcome can be implemented with arbitrary precision by using a sufficiently large number of populations.
@@ 4 Partial Implementation
It is immediate that any BWE outcome, for any possible information structure (direct or indirect), is in the set of BCWE. This simple observation explicitly illustrates how to construct some BCWE: choose an information structure $\left(\boldsymbol{\gamma},\mathcal{T},\pi\right)$ for disclosing information, and choose a BWE of the basic game extended with this information structure.
The next theorem establishes the other direction of the revelation principle (Myerson,, 1982, Bergemann and Morris,, 2016), at least approximately. Specifically, it shows that if $\mu$ is a BCWE with finite support and rational flows, then there exists a direct information structure such that $\mu$ is implemented as an obedient BWE outcome. If the support of $\mu$ is not finite or if some flows in the support are not rational, then $\mu$ can be approximately implemented as a Bayesian Wardrop $\varepsilon$-equilibrium outcome for every $\varepsilon>0$.
@@
Let $\mu$ be a BCWE and $\psi\colon\mathcal{Y}\times\Theta\to\mathbb{R}$ be a continuous function. For all $\varepsilon>0$, there exists a direct information structure $\mathcal{I}$ such that the obedient interim flow profile $\breve{\boldsymbol{y}}_{\mathcal{I}}$ is a Bayesian Wardrop $\varepsilon$-equilibrium of $(\Gamma,\mathcal{I})$ and
@@
In Theorem˜4.1, the function $\psi$ represents any possible designer welfare function, and Eq.˜4.1 means that for every state, the interim expected payoff of the designer under outcome $\breve{\mu}_{\mathcal{I}}$ is $\varepsilon$-close to the interim expected payoff under the BCWE $\mu$. The proof of the theorem also shows that, for every $L>0$, the same $\mathcal{I}$ can be chosen for all $L$-Lipschitz functions $\psi$.
Proof of Theorem˜4.1.
Koessler et al., (2025, Lemma 2) prove that BCWE with finite support are dense within the set of Bayes correlated Wardrop equilibria. So, it is enough to prove the result for $\mu$ with finite support. Fix $\varepsilon>0$ and a BCWE $\mu$ with finite support $\mathcal{Y}^{\ast}\subset\mathcal{Y}$. Let $\mathcal{A}^{*}$ be the set of actions $a$ actually used under $\mu$, i.e., such that
$\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}p(\theta)\mu(\boldsymbol{y}\mid\theta)y_{a}>0.$ (4.2)
Let then
$\varepsilon_{0}=\min\Big\{\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}p(\theta)\mu(\boldsymbol{y}\mid\theta)y_{a}\text{ s.t. }a\in\mathcal{A}^{*}\Big\}>0.$ (4.3)
Fix an arbitrary $\eta>0$ and approximate flows in $\mathcal{Y}^{\ast}$ by vectors with rational components as follows. There exists $K\in\mathbb{N}_{+}$, such that for every $\boldsymbol{y}\in\mathcal{Y}^{\ast}$ and every $a\in\mathcal{A}^{*}$, there exists $N(y_{a})\in\mathbb{N}$, such that,
$\sum_{a\in\mathcal{A}^{*}}N(y_{a})=K\quad\text{and}\quad\forall a\in\mathcal{A}^{*},\left\lvert\frac{N(y_{a})}{K}-y_{a}\right\rvert\leq\eta.$ (4.4)
For each $\boldsymbol{y}\in\mathcal{Y}^{\ast}$, let $\widetilde{\boldsymbol{y}}$ denote the flow profile $(\widetilde{y}_{a})_{a\in\mathcal{A}^{*}}$ with $\widetilde{y}_{a}=N(y_{a})/K$.
We construct a direct information structure $\mathcal{I}$ as follows. There are $|\mathcal{K}|=K$ populations and each population $k$ has the same size $\gamma^{k}=1/K$. For each population $k$, the set of types is $\mathcal{T}^{k}=\mathcal{A}^{*}$. Conditionally on state $\theta$, a flow profile $\boldsymbol{y}\in\mathcal{Y}^{\ast}$ is drawn with probability $\mu(\boldsymbol{y}\mid\theta)$, then action $a$ is recommended to a subset of $N(y_{a})$ populations chosen uniformly, so that, conditionally on $\boldsymbol{y}$ and $\theta$, the probability that population $k$ is recommended $a$ is $\widetilde{y}_{a}$. We let $\breve{\mu}_{\mathcal{I}}$ be the induced obedient outcome which puts probability $\mu(\boldsymbol{y}\mid\theta)$ on $\widetilde{\boldsymbol{y}}$ in state $\theta$.
For each population $k$, the conditional expected payoff of playing $b$ when $a$ is recommended is
$\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}\frac{p(\theta)\mu(\boldsymbol{y}\mid\theta)\widetilde{y}_{a}}{\operatorname{\mathsf{Q}}(a)}u_{b}(\widetilde{\boldsymbol{y}},\theta),$ (4.5)
@@
$\operatorname{\mathsf{Q}}(a)=\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}p(\theta)\mu(\boldsymbol{y}\mid\theta)\widetilde{y}_{a}.$ (4.6)
For each pair of actions $(a,b)$, the mapping $\boldsymbol{y}\mapsto y_{a}u_{b}(\boldsymbol{y},\theta)$ is uniformly continuous on the compact set $\mathcal{Y}$. Thus, there exists a common modulus of continuity $\omega(\,\cdot\,)$ with $\lim_{\eta\searrow 0}\omega(\eta)=0$ such that, for all $\theta\in\Theta$, for all $(a,b)\in\mathcal{A}\times\mathcal{A}$, for all $(\boldsymbol{y},\boldsymbol{z})\in\mathcal{Y}\times\mathcal{Y}$,
[…]
@@
$\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}p(\theta)\mu(\boldsymbol{y}\mid\theta)y_{a}u_{a}(\boldsymbol{y},\theta)\geq\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}p(\theta)\mu(\boldsymbol{y}\mid\theta)y_{a}u_{b}(\boldsymbol{y},\theta).$ (4.8)
From Eqs.˜4.8, 4.4 and 4.7 it follows that
$\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}p(\theta)\mu(\boldsymbol{y}\mid\theta)\widetilde{y}_{a}u_{a}(\widetilde{\boldsymbol{y}},\theta)\geq\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}p(\theta)\mu(\boldsymbol{y}\mid\theta)\widetilde{y}_{a}u_{b}(\widetilde{\boldsymbol{y}},\theta)-2\omega(\eta).$ (4.9)
Thus
$\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}\frac{p(\theta)\mu(\boldsymbol{y}\mid\theta)\widetilde{y}_{a}}{\operatorname{\mathsf{Q}}(a)}u_{a}(\widetilde{\boldsymbol{y}},\theta)\geq\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}\frac{p(\theta)\mu(\boldsymbol{y}\mid\theta)\widetilde{y}_{a}}{\operatorname{\mathsf{Q}}(a)}u_{b}(\widetilde{\boldsymbol{y}},\theta)-\frac{2\omega(\eta)}{\operatorname{\mathsf{Q}}(a)}.$ (4.10)
From Eq.˜4.4 we have
$\operatorname{\mathsf{Q}}(a)=\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}p(\theta)\mu(\boldsymbol{y}\mid\theta)\widetilde{y}_{a}\geq\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}p(\theta)\mu(\boldsymbol{y}\mid\theta)(y_{a}-\eta)\geq\varepsilon_{0}-\eta\geq\frac{\varepsilon_{0}}{2},$ (4.11)
@@
Now, take a continuous $\psi\colon\mathcal{Y}\times\Theta\to\mathbb{R}$. Since $\mathcal{Y}\times\Theta$ is compact, $\psi$ is uniformly continuous and admits a modulus of continuity $\omega_{\psi}$ such that for all $\theta\in\Theta$, for all $(\boldsymbol{y},\boldsymbol{z})\in\mathcal{Y}\times\mathcal{Y}$,
[…]
For each $\theta\in\Theta$, we have
$\begin{split}\left\lvert\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}\mu(\boldsymbol{y}\mid\theta)\psi(\widetilde{\boldsymbol{y}},\theta)-\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}\mu(\boldsymbol{y}\mid\theta)\psi(\boldsymbol{y},\theta)\right\rvert&\leq\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}\mu(\boldsymbol{y}\mid\theta)\left\lvert\psi(\widetilde{\boldsymbol{y}},\theta)-\psi(\boldsymbol{y},\theta)\right\rvert\\ &\leq\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}\mu(\boldsymbol{y}\mid\theta)\omega_{\psi}(\eta)\\ &\leq\omega_{\psi}(\eta),\end{split}$ (4.14)
@@
If flows in $\mathcal{Y}^{\ast}$ have rational coefficients with common denominator $K$, then we can choose $\widetilde{\boldsymbol{y}}=\boldsymbol{y}$ and $\eta=0$, and we have a Bayesian Wardrop 0-equilibrium. Finally, notice that if $\psi$ is $L$-Lipschitz (for the sup norm on flows), then we can choose $\omega_{\psi}(\eta)=L\eta$. Thus, for given $\varepsilon$, the choice of $\eta$ and $\mathcal{I}$ is the same across all such $\psi$. ∎
5 Potential Games and Full Implementation
In this section, we study full implementation of BCWE in games with a concave potential.
5.1 Potential Games